# Scoring Intervals using Non-hierarchical Transformer for Automatic Piano Transcription

### v4

@@ Abstract
The neural semi-Markov Conditional Random Field (semi-CRF) framework has demonstrated promise for event-based piano transcription. In this framework, all events (notes or pedals) are represented as closed intervals tied to specific event types. The neural semi-CRF approach requires an interval scoring matrix that assigns a score for every candidate interval. However, designing an efficient and expressive architecture for scoring intervals is not trivial. In this paper, we introduce a simple method for scoring intervals using scaled inner product operations that resemble how attention scoring is done in transformers. We show theoretically that, due to the special structure from encoding the non-overlapping intervals, under a mild condition, the inner product operations are expressive enough to represent an ideal scoring matrix that can yield the correct transcription result. We then demonstrate that an encoder-only non-hierarchical transformer backbone, operating only on a low-time-resolution feature map, is capable of transcribing piano notes and pedals with high accuracy and time precision. The experiment shows that our approach achieves the new state-of-the-art performance across all subtasks in terms of the F1 measure on the Maestro dataset.
@@ 1 Introduction
Automatic Music Transcription (AMT) transforms the audio signal of music performances into symbolic representations [1]. In this work, we focus on transcribing piano performance audio into its piano roll representation. [^1] Preprint, Under-review. Code will be available upon publication.
@@
Our study focuses on the neural semi-Markov Conditional Random Field (semi-CRF) framework [2] from the second category, which directly models each music event (note or pedal) as a closed interval associated with a specific event type. The approach employs a neural network to score interval candidates and uses dynamic programming to decode non-overlapping intervals. This framework eliminates the need for separate keypoint detection and assembly steps in the first category but outputs the events (intervals) in a single stage. Compared to other methods in the second category, e.g. [6, 7], it does not need hand-crafted state definitions and state transitions. Additionally, it benefits from optimal decoding in a non-autoregressive fashion as opposed to the slow autoregressive and suboptimal decoding in seq2seq methods (the third category).
Although the framework adopted from [2] and the technique proposed in this paper are generally applicable to AMT of other instruments and sound event detection for general audio, we limit our scope to piano transcription for the availability of high-quality nonsynthetic datasets, such as Maestro [9], MAPS [10] and SMD [11].
This paper builds upon, simplifies, and improves the neural semi-CRF framework [2] for piano transcription. Our major contributions are as follows. First, we replace the original scoring module that assigns a score for every possible interval with a simpler and more efficient pairwise inner product operation. Specifically, we prove that due to the special structure of encoding non-overlapping intervals, under a mild condition, the inner product operation is expressive enough to represent an ideal scoring matrix that can yield the correct transcription decoding. Second, inspired by the resemblance between the proposed inner product operation and the attention mechanism in the transformer [12], we use the transformer architecture to produce the interval representation for inner product scoring. We demonstrate that an encoder-only non-hierarchical transformer backbone, operating only on a low-time-resolution feature map, is capable of transcribing notes with high accuracy and time precision. Third, we compare our method against state-of-the-art piano transcription algorithms on the Maestro v3 dataset, showing that our method establishes the new state of the art across all subtasks in terms of the F1 score.
@@ 2.1 Neural Semi-CRF for Piano Transcription
Here, $\operatorname*{arg\,max}$ in Eq. 2, and the summation in the second term in Eq. 3 are over all possible set of non-overlapping intervals. We refer the readers to [2] for algorithmic details.
@@
2.2 Vision Transformer
In recent years, the Vision Transformer (ViT) [13] marks a significant architectural shift in computer vision, offering an effective alternative to traditional convolutional neural network (CNN) models. ViT segments an image into fixed-size patches, which are processed by transformer layers [12], and has proven successful in various computer vision tasks. The architecture has also been adapted for audio analysis [14, 15] and has achieved success. Our neural architecture design resembles the plain ViT [16, 17], which applies the transformer encoder only on the low-resolution feature map for object detection.
@@ 3.2 Interval Scoring with Inner Product
Given that no two non-zero entries in ${\bm{S}}^{\prime}_{{\mathcal{Y}}}$ share a row or column (as per the definition of set of non-overlapping intervals), and there are $M$ non-zero entries, the rank of ${\bm{S}}^{\prime}_{{\mathcal{Y}}}$ is $M$. Since there are at most $T-1$ non-overlapping intervals across $T$ frames, we have $M\leq T-1$, and the number of nonzero entries in $S^{\prime}_{{\mathcal{Y}}}$ is smaller than or equal to $T-1$, as a result, $-\epsilon{\bm{1}}$ ($T$ non-zeros) cannot be represented by a linear combination of other nonzero columns in ${\bm{S}}^{\prime}_{{\mathcal{Y}}}$, therefore ${rank({\bm{S}}_{{\mathcal{Y}}})=rank({\bm{S}}^{\prime}_{{\mathcal{Y}}})+1=M+1}$. ∎
@@
From Theorem 3.2, by applying a scaling factor [^4] Note that applying a length-dependent scaling on the ideal scoring matrix does not change the decoded result. and reintegrating diagonal terms, we can recover Eq. 5.
@@
    Consider a linear upsampling operator $u_{c}$, which is a special case of a 1-d transposed convolutional layer. It works by dividing each step of a vector sequence into $c$ equal parts when the sequence is upsampled $c$ times. Suppose we want to represent ${\bm{Q}}_{\mathcal{Y}}$ and ${\bm{K}}_{{\mathcal{Y}}}$ using low-resolution 1-d vector sequences: ${\bm{Q}}^{\prime}_{{\mathcal{Y}}}=[{\bm{q}}^{\prime}_{0},\ldots,{\bm{q}}^{\prime}_{T^{\prime}-1}]$ and ${\bm{K}}^{\prime}_{{\mathcal{Y}}}=[{\bm{k}}^{\prime}_{0},\ldots,{\bm{k}}^{\prime}_{T^{\prime}-1}]$ where $T^{\prime}<T$, and this representation is achieved by applying $u_{c}$ to ${\bm{Q}}^{\prime}_{{\mathcal{Y}}}$ and ${\bm{K}}^{\prime}_{{\mathcal{Y}}}$, resulting in ${\bm{Q}}_{\mathcal{Y}}=u_{c}({\bm{Q}}^{\prime}_{\mathcal{Y}})$, and ${\bm{K}}_{\mathcal{Y}}=u_{c}({\bm{K}}^{\prime}_{\mathcal{Y}})$, where $c=T/T^{\prime}$ represents the upsampling factor. For this representation to be valid, the vector dimension $D^{\prime}$ for the low-dimensional sequence, i.e., ${\bm{q}}^{\prime}_{i}$ and ${\bm{k}}^{\prime}_{i}$ should exceed $c|{\mathcal{Y}}|$.
These observations highlight that the dimensionality requirement depends solely on the count of intervals in $\mathcal{Y}$ and the downsampling (upsampling) factor $c=T/T^{\prime}$ along the time axis. This analysis reveals sufficient conditions to guarantee the expressiveness of the inner product interval scoring method.
@@ 3.3 Comparison with Attention Mechanism
Comparing the neural semi-CRF with the inner product scoring to the attention mechanism reveals interesting parallels. The original score module, as in [2], resembles an additive attention mechanism, as introduced by [18]. However, attention mechanisms based on inner products [19] have become preferred for their simplicity and computational efficiency. Similarly, the proposed inner product scoring for neural semi-CRFs efficiently scores intervals. However, in contrast to attention mechanisms that score sequence positions and normalize posteriors for each position, neural semi-CRFs score intervals and normalize posteriors globally over sets of non-overlapping intervals.
@@ 4 Proposed System
Figure 1: Overview of the proposed system.
Figure 1 summarizes the proposed system. The input is an oversampled log-mel spectrogram, as in [2]. The spectrogram is downsampled using strided convolutional layers, followed by the addition of spatial position embeddings (Section 4.2). Event tracks for all event types (notes and pedals) are initialized with their own spatial position embeddings and concatenated with the downsampled spectrogram representations. The concatenated features are processed by a transformer encoder. Subsequently, only the event track embeddings are upsampled using one 1D transposed convolutional layer. The upsampled event tracks are used for inner product interval scoring (Eq. 5) to generate interval scoring matrices, which are then fed to the neural semi-CRF layer for log-likelihood calculation or inference.
@@ 4.1 Rethinking Downsampling
Existing studies on Vision Transformers (ViTs) demonstrate the effectiveness of using highly downsampled, low-resolution feature maps even for tasks requiring dense predictions, e.g., [16]. However, state-of-the-art (SOTA) piano transcription systems, including [2, 5, 4, 8], retain full resolution along the time axis. These approaches preserve the temporal detail of the input frames, but at the cost of increased training time and reduced model scalability.
@@
To upsample event tracks to the original temporal resolution of frames, we utilize a single transposed 1D convolutional layer. We found that this simple upsampling layer efficiently prepares representations for inner product scoring at the desired resolution.
@@ 4.2 Transformer Encoder Architecture
where ${\bm{W}}_{r}$ is a learnable matrix $\mathbb{R}^{B\times C}$, initialized from $\mathcal{N}(0,\gamma^{-2})$; $B$ is the dimension for the Fourier features; $\gamma$ is a hyperparameter; ${\bm{b}}\in\mathbb{R}^{B}$ is the learnable bias term, initialized from $\mathcal{U}(-\pi,+\pi)$; $g:\mathbb{R}^{B}\to\mathbb{R}^{E}$ is a two-layered perceptron. This position embedding functions like an MLP that takes coordinates as input, with the first nonlinearity being a scaled cosine function.
@@
Figure 2(a) illustrates the basic transformer block. This block first applies RMSNorm[22] before the self-attention and feed-forward layers. To enhance training stability, we use ReZero [23] which applies a learnable scaling factor $\lambda$, initially set to $0.01$, before adding to the skip connection. As in Figure 2(b), for reducing computational cost, we alternate attention within each transformer block along the time and frequency/eventType axes; similar ideas are often used for efficient transformer architectures [24, 15].
@@ 5.1 Dataset
The MAPS dataset includes both synthesized and real piano recordings, with the real recordings captured by MIDI playback on Yamaha Disklavier. We evaluate our model on the Disklavier subset (ENSTDkAm/MUS and ENSTDkCl/MUS) of the MAPS dataset, which consists of 60 recordings and is commonly used for cross-dataset evaluation. However, we discovered systematic alignment issues in the ground truth annotations for both notes and pedals, affecting both onset and offset locations. Onset alignment issues have been previously reported [26] but are not widely known in the community. 666 A piece-dependent onset latency around 15 ms has been previously discussed in [26]. Due to the electro-mechanical playback mechanism, this latency could also be note/pedal dependent. Offset deviation (up to approximately 70 ms) appears more complex and may be influenced by pedal-/note-dependent mechanical latency.
@@
Similar to MAPS Disklavier portion, the SMD dataset was created by playing back MIDI files on a Yamaha Disklavier. The dataset contains 50 recordings. We found that the onset annotations in SMD are better aligned compared to MAPS, but the offset annotations suffer from similar biases.
@@ 5.2 Model Specification
A brief summary of the model specifications is listed in Table 1. Training takes about 6 days on 2 NVIDIA RTX 4090.
@@
Upsampling 1d. transposed conv, out: 256, kernel size:8, stride:8
@@ 5.4 Results
Our results on the Maestro v3 test set are presented in Table 2. The proposed model achieves state-of-the-art performance across all metrics in terms of f1 score, surpassing previous methods by a significant margin.
We also report results for soft pedal transcription, a task that, to our knowledge, has not been previously explored. However, the low event-level metrics for soft pedals suggest that accurately determining their onset and offset times is more challenging compared to notes and sustain pedals.
@@
All methods exhibit low activation-level F1 scores on both MAPS and SMD. Using the onset-corrected annotation (Cogliati) on MAPS degrades the activation-level F1 score due to the offset inaccuracy. In fact, the Cogliati annotation achieves similar or lower activation-level F1 scores compared to all listed methods when evaluated against the original annotation.
@@
These observations highlight the need for caution when evaluating models on datasets created using electromechanical playback devices like the Yamaha Disklavier. Despite these complications, our proposed system, with or without data augmentation [^7] Data augmentation: pitch shifting ±$\pm 20$ cents, adding noise from [30], applying randomized 8 band EQ and impulse response from [31]. , achieves the highest note onset F1 score among the compared methods on both SMD and MAPS with Ad hoc/Cogliati correction.
@@
Between Ground Truth

### v5

# Scoring Time Intervals using Non-Hierarchical Transformer for Automatic Piano Transcription

@@ Abstract
The neural semi-Markov Conditional Random Field (semi-CRF) framework has demonstrated promise for event-based piano transcription. In this framework, all events (notes or pedals) are represented as closed time intervals tied to specific event types. The neural semi-CRF approach requires an interval scoring matrix that assigns a score for every candidate interval. However, designing an efficient and expressive architecture for scoring intervals is not trivial. This paper introduces a simple method for scoring intervals using scaled inner product operations that resemble how attention scoring is done in transformers. We show theoretically that, due to the special structure from encoding the non-overlapping intervals, under a mild condition, the inner product operations are expressive enough to represent an ideal scoring matrix that can yield the correct transcription result. We then demonstrate that an encoder-only structured non-hierarchical transformer backbone, operating only on a low-time-resolution feature map, is capable of transcribing piano notes and pedals with high accuracy and time precision. The experiment shows that our approach achieves the new state-of-the-art performance across all subtasks in terms of the F1 measure on the Maestro dataset.
@@ 1 Introduction
Our study focuses on the neural semi-Markov Conditional Random Field (semi-CRF) framework [2] from the second category, which directly models each music event (note or pedal) as a closed time interval associated with a specific event type. The approach employs a neural network to score interval candidates and uses dynamic programming to decode non-overlapping intervals. This framework eliminates the need for separate keypoint detection and assembly steps in the first category but outputs the events (intervals) in a single stage. Compared to other methods in the second category, e.g. [6, 7], it does not need hand-crafted state definitions and state transitions. Additionally, it benefits from optimal decoding in a non-autoregressive fashion as opposed to the slow autoregressive and suboptimal decoding in Seq2Seq methods (the third category).
This paper builds upon, simplifies, and improves the neural semi-CRF framework [2] for piano transcription. Our major contributions are as follows. First, we replace the original scoring module that assigns a score for every possible interval with a simpler and more efficient pairwise inner product operation. Specifically, we prove that due to the special structure of encoding non-overlapping intervals, under a mild condition, the inner product operation is expressive enough to represent an ideal scoring matrix that can yield the correct transcription decoding. Second, inspired by the resemblance between the proposed inner product operation and the attention mechanism in the transformer [9], we use the transformer architecture to produce the interval representation for inner product scoring. We demonstrate that an encoder-only non-hierarchical transformer backbone, operating only on a low-time-resolution feature map, is capable of transcribing notes with high accuracy and time precision. Third, we compare our method against state-of-the-art piano transcription systems on the Maestro v3 dataset, showing that our method establishes the new state of the art across all subtasks in terms of the F1 score.
@@ 2.1 Neural Semi-CRF for Piano Transcription
Here, $\operatorname*{arg\,max}$ in Eq. 2, and the summation in the second term in Eq. 3 are over all possible sets of non-overlapping intervals. We refer the readers to [2] for algorithmic details.
@@
2.2 Vision Transformer and YOLOS
The Vision Transformer (ViT) [10] introduced a significant shift in computer vision, offering an alternative to traditional CNN models. ViT processes images as sequences of fixed-size patches using transformer layers [9], proving successful across various tasks. For end-to-end object detection, YOLOS [11] demonstrated a minimal, non-hierarchical encoder-only design that appends [DET] tokens (representing object slots) directly to image patch tokens as input to the transformer encoder. Our architecture adopts a similar encoder-only design for event-based music transcription.
@@ 3.2 Interval Scoring with Inner Product
Given that no two non-zero entries in ${\bm{S}}^{\prime}_{{\mathcal{Y}}}$ share a row or column (as per the definition of set of non-overlapping intervals), and there are $M$ non-zero entries, the rank of ${\bm{S}}^{\prime}_{{\mathcal{Y}}}$ is $M$. Since there are at most $T-1$ non-overlapping intervals across $T$ frames, we have $M\leq T-1$, and the number of nonzero entries in $S^{\prime}_{{\mathcal{Y}}}$ is smaller than or equal to $T-1$. As a result, $-\epsilon{\bm{1}}$ ($T$ non-zeros) cannot be represented by a linear combination of other nonzero columns in ${\bm{S}}^{\prime}_{{\mathcal{Y}}}$, therefore ${rank({\bm{S}}_{{\mathcal{Y}}})=rank({\bm{S}}^{\prime}_{{\mathcal{Y}}})+1=M+1}$. ∎
@@
    Consider a linear upsampling operator $u_{c}$, which is a special case of a 1-d transposed convolutional layer. It works by dividing each step of a vector sequence into $c$ equal parts when the sequence is upsampled $c$ times. Suppose we want to represent ${\bm{Q}}_{\mathcal{Y}}$ and ${\bm{K}}_{{\mathcal{Y}}}$ using low-resolution 1-d vector sequences: ${\bm{Q}}^{\prime}_{{\mathcal{Y}}}=[{\bm{q}}^{\prime}_{0},\ldots,{\bm{q}}^{\prime}_{T^{\prime}-1}]$ and ${\bm{K}}^{\prime}_{{\mathcal{Y}}}=[{\bm{k}}^{\prime}_{0},\ldots,{\bm{k}}^{\prime}_{T^{\prime}-1}]$ where $T^{\prime}<T$, and this representation is achieved by applying $u_{c}$ to ${\bm{Q}}^{\prime}_{{\mathcal{Y}}}$ and ${\bm{K}}^{\prime}_{{\mathcal{Y}}}$, resulting in ${\bm{Q}}_{\mathcal{Y}}=u_{c}({\bm{Q}}^{\prime}_{\mathcal{Y}})$, and ${\bm{K}}_{\mathcal{Y}}=u_{c}({\bm{K}}^{\prime}_{\mathcal{Y}})$, where $c=T/T^{\prime}$ represents the upsampling factor. For this representation to be valid, the vector dimension $D^{\prime}$ for the low-resolution sequence, i.e., ${\bm{q}}^{\prime}_{i}$ and ${\bm{k}}^{\prime}_{i}$ should exceed $c|{\mathcal{Y}}|$.
These observations highlight that the dimensionality requirement depends solely on the count of intervals in $\mathcal{Y}$ and the downsampling (upsampling) factor $c=T/T^{\prime}$ along the time axis. This analysis reveals sufficient conditions to guarantee the expressiveness of the inner product interval scoring method. From Theorem 3.2, by applying a scaling factor [^4] Note that applying a length-dependent scaling on the ideal scoring matrix does not change the decoded result. and reintegrating diagonal terms, we can recover Eq. 5.
@@ 3.3 Comparison with Attention Mechanism
Comparing the neural semi-CRF with the inner product scoring to the attention mechanism reveals interesting parallels. The original score module, as in [2], resembles an additive attention mechanism, as introduced by [12]. However, attention mechanisms based on inner products [13] have become preferred for their simplicity and computational efficiency. Similarly, the proposed inner product scoring for neural semi-CRFs efficiently scores intervals. Computationally, semi-CRFs have the same quadratic computational time complexity as attention heads. However, in contrast to attention mechanisms that score sequence positions and normalize posteriors for each position, neural semi-CRFs score intervals and normalize posteriors globally over sets of non-overlapping intervals.
@@ 4 Proposed System
Figure 1: Overview of the proposed system. Inner product scoring follows Eq. 5.
Figure 1 summarizes the proposed system. The input is an oversampled log-mel spectrogram, as in [2]. The spectrogram is downsampled using 2-d strided convolutional layers, followed by the addition of spatial position embeddings (Section 4.2). Event tracks for all event types (notes and pedals) are initialized with their own spatial position embeddings and concatenated with the downsampled spectrogram representations. The concatenated features are processed by a transformer encoder. Subsequently, only the event track embeddings are upsampled using one 1-d transposed convolutional layer. The upsampled event tracks are used for inner product interval scoring (Eq. 5) to generate interval scoring matrices, which are then fed to the neural semi-CRF layer for log-likelihood calculation or inference.
@@ 4.1 Rethinking Downsampling
Existing studies on Vision Transformers (ViTs) demonstrate the effectiveness of a non-hierarchical design that uses highly downsampled, low-resolution feature maps even for tasks requiring dense predictions, e.g., [14], challenging the dominance of hierarchical models like UNET [15]. However, state-of-the-art (SOTA) piano transcription systems, including [2, 5, 4, 8], retain full resolution along the time axis. These approaches preserve the temporal detail of the input frames, but at the cost of increased training time and reduced model scalability.
@@
To upsample event tracks to the original temporal resolution of frames, we utilize a single transposed 1-d convolutional layer. We found that this simple upsampling layer efficiently prepares representations for inner product scoring at the desired resolution.
@@ 4.2 Transformer Encoder Architecture
where ${\bm{W}}_{r}$ is a learnable matrix $\mathbb{R}^{B\times C}$, initialized from $\mathcal{N}(0,\gamma^{-2})$; $B$ is the dimension for the Fourier features; $\gamma$ is a hyperparameter; ${\bm{b}}\in\mathbb{R}^{B}$ is the learnable bias term, initialized from $\mathcal{U}(-\pi,+\pi)$; $g:\mathbb{R}^{B}\to\mathbb{R}^{E}$ is a two-layer perceptron. This position embedding functions like an MLP that takes coordinates as input, with the first nonlinearity being a scaled cosine function.
@@
Figure 2(a) illustrates the basic transformer block. This block first applies RMSNorm[18] before the self-attention and feed-forward layers. To enhance training stability, we use ReZero [19] which applies a learnable scaling factor $\lambda$, initially set to $0.01$, before adding to the skip connection. As in Figure 2(b), for reducing computational cost, we alternate attention within each transformer block along the time and frequency/eventType axes; similar ideas are often used for efficient transformer architectures [20, 21, 22].
@@ 5.1 Dataset
The MAPS dataset includes both synthesized and real piano recordings, with the real recordings captured by MIDI playback on Yamaha Disklavier. We evaluate our model on the Disklavier subset (ENSTDkAm/MUS and ENSTDkCl/MUS) of the MAPS dataset, which consists of 60 recordings and is commonly used for cross-dataset evaluation. However, we discovered systematic alignment issues in the ground-truth annotations for both notes and pedals, affecting both onset and offset locations. Onset alignment issues have been previously reported in [26] but are not widely known in the community666 A piece-dependent onset latency around 15 ms has been previously discussed in [26]. Due to the electro-mechanical playback mechanism, this latency could also be note/pedal dependent. Offset deviation (up to approximately 70 ms) appears more complex and may be influenced by pedal-/note-dependent mechanical latency or undocumented specific piano model’s response to non-binary pedal values. .
@@
Similar to Maestro dataset, the SMD dataset was created by recording human performance on a Yamaha Disklavier. The dataset contains 50 recordings. We found that the onset annotations in SMD are better aligned compared to MAPS, but the offset annotations suffer from similar biases.
@@ 5.2 Model Specification
The key model specifications are summarized in Table 1. Training takes about 6 days on 2 NVIDIA RTX 4090.
@@
Upsampling 1d. transposed conv, out: 128, kernel size:8, stride:8
@@ 5.4 Results
Our results on the Maestro v3 test set are presented in Table 2. The proposed model achieves state-of-the-art performance across all metrics in terms of f1 score, surpassing previous methods by a significant margin. We also report results for soft pedal transcription which has not been previously explored. The low event-level metrics suggest that accurately determining soft pedal onset and offset times is more challenging than for notes and sustain pedals. We conjecture this is because soft pedals are typically engaged for longer durations and appear significantly less frequently in the dataset than sustain pedals.
@@
Effect of omitting incomplete events.
We found that omitting steps of handling incomplete events at segment boundaries (Section 4.3) only cause noticeable performance impact for pedals, particularly the soft pedal (Table 2). This can be explained by the fact that pedal events, especially soft pedals, can often exceed the segment length, while notes are normally shorter than the segment length we choose.
@@
All methods exhibit low activation-level F1 scores on both MAPS and SMD. Using the onset-corrected annotation (Cogliati) on MAPS increases the onset F1 score but degrades the activation-level F1 score due to the uncorrected offset biases. In fact, the Cogliati annotation achieves similar or lower activation-level F1 scores compared to all listed methods when evaluated against the original annotation.
@@
These observations highlight the need for caution when evaluating models on datasets created using mechanisms that may involve systematic biases, e.g., electromechanical playback. Despite these complications, our proposed system, with or without data augmentation [^7] Data augmentation: pitch shifting ±$\pm 20$ cents, adding noise from [31], applying randomized 8 band EQ and impulse response from [32]. , achieves the highest note onset F1 score among the compared methods on both SMD and MAPS with Ad hoc/Cogliati correction.
@@
00footnotetext: Previous SOTA for notes. Use their provided code and pretrained weights. Recomputed from transcribed MIDIs. 00footnotetext: Previous SOTA for sustain pedals. Their released code indicates a 200 ms onset tolerance for pedal evaluation, contrary to the reported 50 ms in their paper. Here, we use a 50 ms onset tolerance, which explains the large discrepancy between the numbers here and their reported results.
@@
Between Ground Truths